Merrimac Brewing company's total assets equal $18 million. The book value of Merrimac's equity is $6 million. Excess cash is $200,000. The market value of Merrimac's equity is $10 million. Its Debt to Enterprise Value ratio is .5. What is Merrimac's Debt Ratio?
Answer #1

Total Assets = $ 18 M

Equity = $ 6 Million

Total Debt = Total Assets - Total Equity = 18 - 6 = 12 M

Debt ratio = Debt/ Assets = 12/18 = 66.67%